Hoe many sulfer atoms are there in 25g of AL2S3?


1
Expert's answer
2021-07-15T02:40:44-0400

Generally, the number of atoms = (mass/molar mass)*Avogadro’s number

First, find the total moles in AI2S3= mass/molar

Molar mass of Al2S3 =150.158g/mol

Mass=25g

Thus, number of moles= 25 g/150.158g/mol=0.167 moles

Since there are three atoms of S3 in Al2S3 the moles of S3 = 3*0.167=0.50 moles

When using Avogadro’s constant, 1 mole = 6.02* 10^23 atoms

Therefore, the number of S atoms in Al2S3= 0.50 * 6.02 x10^23= 3.01*10^23